CREAM Flight 2008
CREAM-IV at McMurdo, Nov. 20th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 20th of November at McMurdo.
- Not so much wind outside the building and the TDRSS antenna located outside did not have any problems during the End-to-End test this afternoon.
Weather briefing shows no significant changes in the stratosphere.
CREAM-IV at McMurdo, Nov. 19th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 19th of November at McMurdo.
- We still have a windy weather expected to coming on Thursday (Wed. in US) even though it would not be a storm.
- Will prepare the End-to-End test tomorrow afternoon as planned.
- TCD PXL PM#1 ADC0 problem solved by replacing the concentrator box.
CREAM-IV at McMurdo, Nov. 18th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 18th of November at McMurdo.
- In the weather briefing, nothing much changed yet. We still have a windy weather expected to coming on Thursday (Wed. in US) when we plan to do the End-to-End test with the ROCC/ESC/SOC. Will prepare the test as planned and change it if weather becomes bad.
- The pattern of stratosphere is currently about 2 days later than 2007 but still expect to have the vortex formed around 10-15th of December.
- The instrument CPT is complete. The calorimeter, t/b SCD, Chercam, TCD tests are done today. Everything works fine except the following items.
CREAM-IV at McMurdo, Nov. 17th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 17th of November at McMurdo.
- In the weather briefing, a bad weather is expected to coming on Thursday (Wed. in US) when we plan to do the End-to-End test with the ROCC/ESC/SOC. Will prepare the test as planned and change it if weather becomes bad.
- The pattern of stratosphere is similar to 2007 and still expect to have the vortex formed around 10-15th of December.
CREAM-IV at McMurdo, Nov. 16th 2008
- Details
Dear,
Here is a summary of CREAM-IV flight preparation status on the 16th of November at McMurdo..
- The SFC2 problem was debugged in the whole afternoon and resolved by replacing the USB cable.
- NSF head office had a visit to the LDB facility and also to our CREAM integration building.



